# Attention-based Class-Conditioned Alignment for Multi-Source Domain Adaptation of Object Detectors

### v3

@@ Abstract
Domain adaptation methods for object detection (OD) strive to mitigate the impact of distribution shifts by promoting feature alignment across source and target domains. Multi-source domain adaptation (MSDA) allows leveraging multiple annotated source datasets and unlabeled target data to improve the accuracy and robustness of the detection model. Most state-of-the-art MSDA methods for OD perform feature alignment in a class-agnostic manner. This is challenging since the objects have unique modal information due to variations in object appearance across domains. A recent prototype-based approach proposed a class-wise alignment, yet it suffers from error accumulation due to noisy pseudo-labels that can negatively affect adaptation with imbalanced data. To overcome these limitations, we propose an attention-based class-conditioned alignment method for MSDA that aligns instances of each object category across domains. In particular, an attention module coupled with an adversarial domain classifier allows learning domain-invariant and class-specific instance representations. Experimental results on multiple benchmarking MSDA datasets indicate that our method outperforms the state-of-the-art methods and is robust to class imbalance using a conceptually simple class-conditioning method. Our code is available here.
@@ 1 Introduction
Given the recent advancements in deep learning, object detectors have demonstrated impressive results on many benchmark problems ranging from closed-world fixed class settings to open-world scenarios [5, 43, 32, 20]. However, their performance typically degrades when there is a domain shift between the training (source) and test (target) data distributions. To address this challenge, several unsupervised domain adaptation (UDA) methods [17, 27, 4, 10] have been proposed to improve OD performance in the presence of a domain shift. The conventional UDA methods assume that the source data comes from a single distribution. In real-world scenarios, multiple source datasets may be available for adaptation with considerable domain shifts among these datasets due to variations in capture conditions, like weather, geographic location, and illumination. This setting is referred to as Multi-Source Domain Adaptation (MSDA). MSDA is challenging as it requires addressing discrepancies among different sources, along with sources and target discrepancies. However, leveraging source data from multiple distributions results in enhanced adaptation, increased generalization, and improved robustness against domain shifts[23, 40].
@@ 3 Proposed Method
The training framework of our proposed ACIA is shown in Fig. 3. Our model is based on the Mean-Teacher (MT) framework [21]. The components responsible for image-level and instance-level alignment are integrated into the MT framework. They are the multi-source discriminator and attention-based class-conditioned instance-level discriminator. The MT training process is comprised of two stages. First, the student model is trained in a supervised manner on all source domain data, typically called the burn-in stage. After this step, the teacher model weights are initialized as an exact copy of the student model weights. Then, mutual learning of teacher-student begins on both the labeled source datasets and unlabeled target data. In the case of source datasets, the same supervised training is continued. For the target dataset images, pseudo labels generated by the teacher model are used for training. Meanwhile, the multi-source discriminator and class-conditioned instance-level discriminator are trained for domain-invariant feature generation and class-wise alignment, respectively. The weights of the teacher model are the exponential moving average (EMA) of the student model weights. At inference time, only the teacher model is employed. The rest of the section describes all the loss functions for the proposed ACIA framework.
@@
(b) Target Data Training. The target dataset is used at the teacher-student mutual learning stage of the training process. For every image from the target domain, two augmented versions are generated – with strong and weak augmentation. For the strong augmentation, color jittering, grayscale, Gaussian blur, and cutout patches are used. For weak augmentation, image rescaling, and horizontal flips are used. The weak augmented image is fed to the teacher model, and the output is filtered by confidence thresholding [21] to generate pseudo-labels $\widetilde{y}$. The student network then receives the strongly augmented version and its loss is computed using the the same loss function as Eq. 1:
$\mathcal{L}_{\mbox{unsup}}=\sum_{i=1}^{M_{T}}\mathcal{L}_{\mbox{cls}}(x^{T}_{i},\widetilde{y}_{i})$ (2)
The regression loss is not used here, as the confidence thresholding is not sufficient to filter the pseudo-labels that are potentially incorrect for bounding box regression [21].
@@
(d) Instance-level Alignment. In our method, the instance-level alignment is performed in a class-wise manner. Each object category has variations in visual attributes (such as appearance, scale, orientation, and other characteristics) across domains. If the variation is significant, it is challenging for the detector to align features across domains, leading to sub-optimal performance [36, 3]. Furthermore, class-wise alignment helps with underrepresented classes by ensuring that the features of these classes are explicitly aligned across domains, regardless of their lower frequency in the data. As shown in Fig. 2, without class-wise alignment, the separation of class features of the same category across domains is high. PMT [3] is the first MSDA approach in OD that implements class-wise alignment using global and local prototypes. The local prototypes ensured within-domain alignment, whereas the global prototypes ensured across-domain alignment. In contrast to using prototypes for each object category and each domain, we use an efficient attention-based module to ensure class-wise instance alignment.
@@ 4.1 Experimental Methodology
(a) Implementation Details. In our experiments, we follow the same settings as [37, 34, 3]. In all experiments, Faster-RCNN [24] is used as the detector, with Imagenet pre-trained VGG-16 backbone. Images are resized such that the shorter side of the image is 600 while maintaining the image ratios, following the ROI-alignment [13] based implementation of Faster-RCNN. We set $\alpha$, $\beta$, and $\gamma$ hyperparameters to 1, 0.1, and 0.3, respectively. The pseudo-label threshold was set to 0.7 for all the experiments. As mentioned in Sec. 3, our training is two-staged. We trained our model for 15 epochs in both the settings. We set the weight smoothing coefficient for the EMA of the teacher model to 0.9996. All experiments were executed on 4 NVIDIA A100 GPUs, with a batch size of 8 and a learning rate of 0.02.
@@ 4.2 Results and Discussion
(a) Cross-time Adaptation. Daytime and Night subsets of the BDD100K dataset are used as the source domain, while the Dusk/Dawn subset of BDD100K is used as the target domain. In this setting, the images are captured at different times. This causes a domain shift due to variations in illumination. The performance of ACIA is compared against other baselines in Tab. 1. Globally, UDA methods outperform the Lower Bound as expected, but there is a slight decrease in performance when the night domain is added. This is because the UDA Blending methods do not consider the domain shift among the source domains, resulting in performance degradation. This emphasizes the importance of considering MSDA methods. ACIA outperforms PMT [3] by 2.6 mAP. Also, our method outperforms Target-Only and All-Combined Upper Bound settings. The poor performance of the Target Only setting is due to fewer samples in the target domain compared to source domains.
@@ D N D+N
Table 1: mAP performance of ACIA and baseline methods for cross-time adaptation on BDD100k, when the 2 sources are daytime (D) and night (N) subsets, and the target is the Dusk/Dawn subset.
@@ 5 Conclusion
6 Supplementary Material
6.1 Datasets
In our paper, we used five datasets to create the multi-source domain adaptation settings. These datasets are listed below and summarized in the Tab. 7:
1. BDD100k - The BDD100K [38] is a large-scale diverse driving dataset. It contains 70,000 training and 10,000 testing images captured across various times, like Daytime, Night, and Dusk/Dawn. This variation makes it a good choice for the DA problem.
2. Cityscapes - Cityscapes [9] is an autonomous driving research dataset with images captured from urban street scenes. It contains 2,975 training and 500 testing images.
3. Kitty - The KITTI [12] dataset is a self-driving dataset that comprises a collection of images and associated sensor data captured from a moving vehicle in urban environments. It consists of 7,481 training images RGB images.
4. MS COCO - MS COCO [19] is one of the most widely used benchmark datasets in computer vision. It is a complex dataset having large scale and appearance variations for the instances. It has around 330,000 images containing 80 object categories.
5. Synscapes - Synscapes [33] is a synthetic autonomous driving dataset that provides more variability for testing our method. It consists of 25,000 training images.
Setting Src.1 # Img. Src.2 # Img. Src.3 # Img. Target # Img.
Cross-Time Day 36,728 Night 27,971 - - Dawn 5,027
Cross-Camera Cityscapes 2,831 Kitty 6,684 - - Day 36,728
Mixed Domain Cityscapes 2,975 COCO 71,745 Synscapes 25,000 Day 36,728
Table 7: Summary of the different MSDA settings used in our study.
6.2 Study on the Class-Embedding layer
In this section, we visualize the information learned by the class-embedding layer. Fig. 5 shows the activation of each class-embedding layer (corresponding to each object category) with ROI-Pooled features that contain the object category. For this, we crop the region of the object category from the image and find activation with each class-embedding layer. After that, we use Softmax to normalize the values. It can be observed from Fig. 5 that each embedding layer is activated most when it matches the corresponding object category. This shows that the class-embedding is successfully learning class information. Also, the under-represented object categories (bike, truck, truck) are highly activated with their corresponding embedding layer. It shows that our attention-based instance-level alignment mechanism is helping in datasets with class imbalance by focusing on under-represented classes.
Figure 5: Heatmap showing the activation of each class embedding with instance features of some object categories. The X-axis represents the list of all the class embedding corresponding to all the object categories, and the Y-axis represents the object category present in the ROI-Pooled feature.
6.3 Increase in number of parameters with source domains
Earlier works in the MSDA for OD [37, 34] relied on learning domain-specific parameters. It rapidly increases training parameters with each source domain, as summarized in Tab. 8. Later, PMT [3] mitigated this by using prototypes instead of domain-specific subnets. However, in PMT the number of parameters slightly increases with each source domain. In contrast, our ACIA doesn’t have any domain-specific parameters, so there is no change in the number of parameters with source domains.
Method Number of source domains
1 2 3 4 5
DMSN 45.994 75.426 104.858 134.290 163.722
TRKP 45.994 59.942 73.890 87.838 101.786
PMT 46.586 46.587 46.588 46.589 46.590
ACIA (ours) 45.659 45.659 45.659 45.659 45.659
Table 8: The number of parameters vs the number of source domains. No parameter growth in our method.
6.4 Additional Experiments on the Mixed Domain Adaptation Setting
In this section, we provide some additional results on the mixed domain adaptation settings. In the main paper, we only showed the results when $C+M$ was employed as the two source domains (because the previous papers followed that setting only). In Tab. 9 we compare our results with PMT [^2] We trained for this setting using the code provided by them when $C+S$ and $M+S$ are employed as the source domain. It can be observed that our method outperforms PMT for both settings. This further shows that our attention-based instance alignment performs better than prototype-based instance alignment with complex domain shifts.
Method C+S M+S
PMT 30.1 34.9
ACIA (ours) 33.7 35.8
Table 9: Additional results on the Mixed Domain Adaptation Setting.
6.5 Effect of Class-Alignment in UDA methods
In this section, we show the effect of class-wise alignment on UDA methods [27, 17]. Tab. 10 shows the results. We studied three different cases: (1) No class alignment - we used their proposed method only. (2) Class-Alignment with target domain - here we incorporated our class alignment component with their method, aligning the source domains and the target domain. (3) Class-Alignment without target domain - our class alignment component with their method, but this time only the source domains are considered. The results clearly show that our class alignment is effective with both methods. Thus, we can conclude that the proposed class-conditional alignment is effective for both UDA and MSDA methods. It can be observed that similar to our method, removing target data for the instance-level alignment further enhances the model performance.
Method No Class Class-Align. Class-Align.
Align. w/ Target w/o Target
Strong-Weak[27] 29.9 33.7 34.2
Adaptive Teacher[17] 34.6 36.8 37.6
Table 10: Importance of Class-alignment in UDA methods. The proposed attention-based class-conditioned aligner is effective for UDA methods as well.
6.6 Architecture of the Discriminator Networks
We use two domain classifiers as the discriminator networks for adversarial training: an image-level domain classifier and an instance-level domain classifier. Their architecture is summarized in Fig. 6. The image-level domain classifier receives its input from the final layer of the backbone network used for feature extraction. This classifier is fully convolutional with a final N+1 class prediction layer (corresponds to the number of source domains plus the target domain). The instance-level classifier receives its input from the attention layer. This classifier consists of only linear layers with a final N-way prediction layer (the target domain is not used here because the GT boxes from the target domain are not available).
Figure 6: Detailed architecture of the networks used for image-level and instance-level domain classifiers. The activation functions used in image-level and instance-level classifiers are leaky ReLU and GELU respectively. Additionally, Layernorm is used in the instance-level domain classifier. (r= no of GT boxes in the image, h,w = height and width of the feature map, FC = Fully connected layer)
6.7 Class-wise AP
We also report the detailed class-wise AP of ACIA on the Cross-Time and Mixed Adaptation settings in Tab. 11 and Tab. 12 respectively. Note that, for the Cross-camera Adaptation settings, there is only one class, so this is not applicable.
In both settings, our method outperforms the others for all classes. The class-wise AP shows improvement both in the majority and minority classes as well - eg: car (majority) and traffic sign (minority) in the cross-time settings. In the mixed adaptation settings, we outperform others for all the classes, in both the case of two and three source domains.
Setting Source Method Bike Bus Car Motor Person Rider Light Sign Train Truck mAP
Lower Bound D Source Only 35.1 51.7 52.6 9.9 31.9 17.8 21.6 36.3 - 47.1 30.4
N 27.9 32.5 49.4 15.0 28.7 21.8 14.0 30.5 - 30.7 25.0
D+N 31.5 46.9 52.9 8.4 29.5 21.6 21.7 34.3 - 42.2 28.9
UDA D+N Strong-Weak[28] 29.7 50.0 52.9 11.0 31.4 21.1 23.3 35.1 - 44.9 29.9
Graph Prototype[4] 31.7 48.8 53.9 20.8 32.0 21.6 20.5 33.7 - 43.1 30.6
Cat. Regularization[35] 25.3 51.3 52.1 17.0 33.4 18.9 20.7 34.8 - 47.9 30.2
UMT[10] 42.3 48.1 56.4 13.5 35.3 26.9 31.1 41.7 - 40.1 33.5
Adaptive Teacher[17] 43.1 48.9 56.9 14.7 36.0 27.1 32.7 43.8 - 42.7 34.6
MSDA D+N MDAN[40] 37.1 29.9 52.8 15.8 35.1 21.6 24.7 38.8 - 20.1 27.6
M3SDA[23] 36.9 25.9 51.9 15.1 35.7 20.5 24.7 38.1 - 15.9 26.5
DMSN[37] 36.5 54.3 55.5 20.4 36.9 27.7 26.4 41.6 - 50.8 35.0
TRKP[34] 48.4 56.3 61.4 22.5 41.5 27.0 41.1 47.9 - 51.9 39.8
PMT[3] 55.3 59.8 67.6 29.9 47.6 32.7 46.3 56.0 - 57.7 45.3
ACIA(Ours) 56.1 61.0 69.2 31.9 51.8 39.8 49.2 59.0 - 61.0 47.9
Upper Bound D+N Target Only 27.2 39.6 51.9 12.7 29.0 15.2 20.0 33.1 - 37.5 26.6
All-Combined 56.4 59.9 67.3 30.8 47.9 33.9 47.2 57.8 - 54.8 45.3
Fine-Tuning 63.3 68.1 72.5 39.3 52.2 37.2 54.1 63.1 - 59.1 50.9
Table 11: Class-wise AP of ACIA compared against the baseline lower bound, UDA, MSDA, and upper bound methods in the cross-time settings. Source domains are daytime (D) and night (N) subsets and the target is always Dusk/Dawn of BDD100K.
Setting Source Method Person Car Rider Truck Motor Bicycle Bike mAP
Lower Bound C Source Only 26.9 44.7 22.1 17.4 17.1 18.8 16.7 23.4
Lower Bound C+M Source Only 35.2 49.5 26.1 25.8 18.9 26.1 26.5 29.7
UDA UMT[10] 30.7 28.0 3.9 11.2 19.2 17.8 18.7 18.5
UDA Adaptive Teacher[17] 31.2 31.7 15.1 16.4 17.1 20.9 27.9 22.9
MSDA TRKP[34] 39.2 53.2 32.4 28.7 25.5 31.1 37.4 35.3
MSDA PMT[3] 41.1 53.5 31.2 31.9 33.7 34.9 44.6 38.7
MSDA ACIA(ours) 43.3 58.1 33.3 35.1 33.7 38.6 45.2 41.0
Upper Bound All-Combined 40.2 60.1 47.1 60.0 29.2 36.3 56.9 47.1
Upper Bound Fine-Tuning 44.1 61.4 49.0 61.1 30.8 39.2 58.8 49.2
Lower Bound C+M+S Source Only 36.6 49.0 22.8 24.9 26.9 28.4 27.7 30.9
UDA UMT[10] 32.7 39.6 6.6 21.2 21.3 25.7 28.5 25.1
UDA Adaptive Teacher[17] 36.3 42.6 19.7 23.4 24.8 27.1 33.2 29.6
MSDA TRKP[34] 40.2 53.9 31.0 30.8 30.4 34.0 39.3 37.1
MSDA PMT[3] 43.3 54.1 32.0 32.6 35.1 36.1 44.8 39.7
MSDA ACIA(ours) 44.9 59.2 33.8 33.5 38.3 39.9 46.5 42.3
Upper Bound All-Combined 41.7 63.9 49.5 58.1 31.6 39.1 53.5 48.2
Upper Bound Fine-Tuning 49.2 63.5 56.1 62.6 35.1 43.7 57.2 52.5
Upper Bound C+M+S Target Only 35.3 53.9 33.2 46.3 25.6 29.3 46.7 38.6
Table 12: Class-wise AP of ACIA compared against the baselines in the mixed adaptation settings. Source domains are Cityscapes(C), MS COCO(M), and Synscapes(S) datasets while the Daytime domain of BDD100K is the target domain.
6.8 More Detection Visualization
In Fig. 7 we present more visualization of the detections on BDD100k cross-time for the three multi-source adaptation approaches presented in the paper: no class-wise adaptation (similar to [37, 34]), prototype-based class-conditional adaptation[3] and our attention-based class-conditional adaptation. From the visualizations, it can be observed that our method is performing better detection compared to the other approaches highlighting the impact of an efficient class-conditioned alignment.
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Figure 7: Comparison of instance-level adaptation detection on BDD100k cross-time setting. (a) MSDA without class-conditioned instance adaptation as in [37, 34]. (b) With the prototype-based class-conditional adaptation [3]. (c) ODs with our ACIA approach.

### v4

@@ Abstract
Domain adaptation methods for object detection (OD) strive to mitigate the impact of distribution shifts by promoting feature alignment across source and target domains. Multi-source domain adaptation (MSDA) allows leveraging multiple annotated source datasets and unlabeled target data to improve the accuracy and robustness of the detection model. Most state-of-the-art MSDA methods for OD perform feature alignment in a class-agnostic manner. This is challenging since the objects have unique modal information due to variations in object appearance across domains. A recent prototype-based approach proposed a class-wise alignment, yet it suffers from error accumulation due to noisy pseudo-labels that can negatively affect adaptation with imbalanced data. To overcome these limitations, we propose an attention-based class-conditioned alignment method for MSDA that aligns instances of each object category across domains. In particular, an attention module coupled with an adversarial domain classifier allows learning domain-invariant and class-specific instance representations. Experimental results [^1] Our code is provided in suppl. materials and will be made public. on multiple benchmarking MSDA datasets indicate that our method outperforms the state-of-the-art methods and is robust to class imbalance using a conceptually simple class-conditioning method.
@@ 1 Introduction
Given the recent advancements in deep learning, object detectors have demonstrated impressive results on many benchmark problems ranging from closed-world fixed class settings to open-world scenarios [5, 43, 32, 20]. However, their performance typically degrades when there is a domain shift between the training (source) and test (target) data distributions. To address this challenge, several unsupervised domain adaptation (UDA) methods [17, 26, 4, 10] have been proposed to improve OD performance in the presence of a domain shift.
The conventional UDA methods assume that the source data comes from a single distribution. In real-world scenarios, multiple source datasets may be available for adaptation with considerable domain shifts among these datasets due to variations in capture conditions, like weather, geographic location, and illumination. This setting is referred to as Multi-Source Domain Adaptation (MSDA). MSDA is challenging as it requires addressing discrepancies among different sources, along with sources and target discrepancies. However, leveraging source data from multiple distributions results in enhanced adaptation, increased generalization, and improved robustness against domain shifts[22, 40].
@@ 3 Proposed Method
The training framework of our proposed ACIA is shown in Fig. 3. Our approach builds on the Mean-Teacher (MT) framework [10], integrating components for both image-level and instance-level alignment. Specifically, these components include a multi-source discriminator for domain-invariant feature learning, and an attention-based instance-level discriminator for class-wise alignment.
The MT training process is comprised of two stages. Firstly, the student model is trained in a supervised manner on all source domain data, typically referred to as the burn-in stage. Following this, the teacher model is initialized as an exact copy of the student model. Mutual learning of the teacher and student begins, utilizing both the labeled source datasets and unlabeled target data. The supervised training continues for the source datasets. Pseudo labels are generated by the teacher model to train on the target dataset.
Meanwhile, the multi-source discriminator and the class-conditioned instance-level discriminator are trained to ensure domain-invariant features and class-wise alignment. Weights of the teacher model are the exponential moving average (EMA) of the student model weights. EMA is used because it is computationally inexpensive. At inference time, only the teacher model is used.
The rest of the section describes all the loss functions for the proposed ACIA framework.
@@
(b) Target Data Training. The target dataset is used at the teacher-student mutual learning stage of the training process. For every image from the target domain, two augmented versions are generated – with strong and weak augmentation. For the strong augmentation, color jittering, grayscale, Gaussian blur, and cutout patches are used. For weak augmentation, image rescaling, and horizontal flips are used. The weak augmented image is fed to the teacher model, and the output is filtered by confidence thresholding [10] to generate pseudo-labels $\widetilde{y}$.
The student network then receives the strongly augmented version and its loss is computed using the the same loss function as Eq. 1:
$\mathcal{L}_{\mbox{unsup}}=\sum_{i=1}^{M_{T}}\mathcal{L}_{\mbox{cls}}(x^{T}_{i},\widetilde{y}_{i})+\mathcal{L}_{\mbox{reg}}(x^{T}_{i},\widetilde{y}_{i})$ (2)
@@
(d) Instance-level Alignment.
In our method, the instance-level alignment is performed in a class-wise manner. Each object category has variations in visual attributes (such as appearance, scale, orientation, and other characteristics) across domains. If the variation is significant, it is challenging for the detector to align features across domains, leading to sub-optimal performance [36, 3]. Furthermore, class-wise alignment helps with underrepresented classes by ensuring that the features of these classes are explicitly aligned across domains, regardless of their lower frequency in the data. As shown in Fig. 2, without class-wise alignment, the separation of class features of the same category across domains is high. PMT [3] is the first MSDA approach in OD that implements class-wise alignment using global and local prototypes. The local prototypes ensured within-domain alignment, whereas the global prototypes ensured across-domain alignment. In contrast to using prototypes for each object category and each domain, we use an efficient attention-based module to ensure class-wise instance alignment.
@@ 4.1 Experimental Methodology
(a) Implementation Details. In our experiments, we follow the same settings as [37, 34, 3]. In all experiments, Faster-RCNN [23] is used as the detector, with Imagenet pre-trained VGG-16 backbone. Images are resized such that the shorter side of the image is 600 while maintaining the image ratios, following the ROI-alignment [13] based implementation of Faster-RCNN. We set $\alpha$, $\beta$, and $\gamma$ hyperparameters to 1, 0.1, and 0.3, respectively. The pseudo-label threshold was set to 0.7 for all the experiments. As mentioned in Sec. 3, our training is two-staged. We trained our model for 15 epochs in both the settings. We set the weight smoothing coefficient for the EMA of the teacher model to 0.9996. All experiments were executed on 4 NVIDIA A100 GPUs, with a batch size of 8 and a learning rate of 0.2.
@@ 4.2 Results and Discussion
(a) Cross-time Adaptation. Daytime and Night subsets of the BDD100K dataset are used as the source domain, while the Dusk/Dawn subset of BDD100K is used as the target domain. In this setting, the images are captured at different times. This causes a domain shift due to variations in illumination. The performance of ACIA is compared against other baselines in Tab. 1. Globally, UDA methods outperform the Lower Bound as expected, but there is a slight decrease in performance when the night domain is added. This is because the UDA Blending methods do not consider the domain shift among the source domains, resulting in performance degradation. This emphasizes the importance of considering MSDA methods. ACIA outperforms PMT [3] by 2.6 mAP. Also, our method outperforms Target-Only and All-Combined Upper Bound settings. The poor performance of the Target Only setting is due to fewer samples in the target domain compared to source domains. In the all combined baseline the model is more biased towards the source datasets. Whereas our method explicitly focuses on improving the performance on the target domain that’s why our result is better.
@@ D N D+N
Table 1: mAP performance of ACIA and baseline methods for cross-time adaptation on BDD100k, when the 2 sources are daytime (D) and night (N) subsets, and the target is the Dusk/Dawn.